OpenStudy (anonymous):

Find the most general antiderivative of the function. G ( t ) = 4 + t + t^2/ sqrt t

OpenStudy (astrophysics):

\[G(t) = 4+t+ \frac{ t^2 }{ \sqrt{t} }\] we can write this as \[G(t) = 4+t+t^{3/2}\] so what do you get when you integrate this expression \[\int\limits (4+t+t^{3/2}) dt\]

OpenStudy (astrophysics):

@Cocheso

OpenStudy (astrophysics):

I wonder if the whole thing is over sqrt t, or just that...

OpenStudy (astrophysics):

\[G(t) = \frac{ 4+t+t^2 }{ \sqrt{t} }\] if so, we can split them up or express it as \[\int\limits (4+t+t^2)t^{-1/2} dt\]